An airplane requires a speed of 80 π‘šπ‘–/β„Žπ‘Ÿ to be airborne. It starts from rest on a runway 1600 𝑓𝑑 long. a) What must be the minimum safe acceleration of the airplane? b) With this acceleration, how many seconds will it take for the plane to acquire its needed speed for take-off?


1
Expert's answer
2021-09-26T08:41:10-0400

Given:

"v=80\:\rm mi/h=35.8\: m/s"

"d=1600\:\rm ft=488\: m"


(a) the minimum safe acceleration of the airplane

"a=\frac{v^2}{2d}=\frac{35.8^2}{2*488}=1.31\:\rm m/s^2=4.30\: ft/s^2"

(b) the time of runway motion

"t=\frac{v}{a}=\frac{35.8}{1.31}=27.3\:\rm s"